OpenStudy (balenator):

A fair die is rolled twice. What is the conditional expectation of the sum of the two outcomes given the first roll shows a 1?

OpenStudy (reemii):

What is the expectation of the result of one single die?

OpenStudy (openstudygirl2):

Balenator i would say the answer is 11/36

OpenStudy (reemii):

E(X+Y | X=1) = E(1+Y) = 1 + E(Y) So what is E(Y) ?

OpenStudy (balenator):

would e(y) be 1/6?

OpenStudy (balenator):

I got 3.5 for E(x+Y), so i would i just do 3.5/E(X=1)?

OpenStudy (phi):

if the first die shows 1 then the sum of the two dice will be 2,3,4,5,6,7 all equally likely and the expected value is the average of those numbers

OpenStudy (balenator):

yes so the expected value would be 3.5

OpenStudy (phi):

I get 27/6 = 9/2 = 4.5

OpenStudy (balenator):

oh sorry i forgot to add 1 to all the them

OpenStudy (balenator):

1/6(1+1)+1/6(1+2)....+1/6(1+6)

OpenStudy (balenator):

would i just divide that by 1/6 since E(x=1) equals 1/6?

OpenStudy (phi):

you do (just as you wrote) 1/6(1+1)+1/6(1+2)....+1/6(1+6) which is the same as doing (2+3+4+5+6+7)/6 we use 1/6 because each sum is equally likely the E(x=1) = 1/6 does not matter, because we are told: given that the first roll is 1 (it does not matter if it's unlikely or it happens all the time... once it happens, we now concentrate on what to expect for the sum of the two dice)

OpenStudy (balenator):

in class we were taught to divide by the given probability, which in this case is (X=1)/ so it would just be 4.5?

OpenStudy (phi):

yes, the answer is 4.5

OpenStudy (phi):

the conditional expectation of the sum of two dice when the first die shows 1 is given, is 4.5

OpenStudy (balenator):

okay thanks